We are evaluating a project that costs $520,000, has a six-year life, and has no salvage value. Assume that depreciation is straight-line to zero over the life of the project. Sales are projected at 65,000 units per year. Price per unit is $45, variable cost per unit is $30, and fixed costs are $840,000 per year. The tax rate is 35 percent, and we require a return of 10 percent on this project. Suppose the projections given for price, quantity, variable costs, and fixed costs are all accurate to within ±10 percent.
Calculate the best-case and worst-case NPV figures. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
a. Best-case $584,621, worst case $ –$285,079 b. Best-case $2,026,176.65, worst case $ –$1,761,593.58 c. Best-case $88,832.77, worst case –$1,510,000 d. None of the above

When I calculated it myself I got :

worst case -$1,761,594.09

best case $2,026,176.14

If you notice it is very close to answer b so I am asking you to calculate it to see if you get a different answer than I did, or am I correct? Thank you
Answer #1

First we calculate the CF for each scenario:

Base Case Best Case Worst Case
Particular Remark Year 0 Year 1-6 Year 1-6 Year 1-6
Units Given $         65,000.00 $       71,500.00 $         58,500.00
SP Given $                 45.00 $                49.50 $                 40.50
Sales Sp x Units $   29,25,000.00 $ 35,39,250.00 $   23,69,250.00
VC Given $                 30.00 $                27.00 $                 33.00
Total VC Vc x Units $   19,50,000.00 $ 19,30,500.00 $   19,30,500.00
Fixed cost Given $     8,40,000.00 $    7,56,000.00 $     9,24,000.00
EBITDA Sale-Total VC-Fixed Cost $     1,35,000.00 $    8,52,750.00 $    -4,85,250.00
Depreciation 520000/6 $         86,666.67 $       86,666.67 $         86,666.67
EBIT EBITDA-Depreciation $         48,333.33 $    7,66,083.33 $    -5,71,916.67
tax 0.35 x EBIT $         16,916.67 $    2,68,129.17 $    -2,00,170.83
EAT EBIT-tax $         31,416.67 $    4,97,954.17 $    -3,71,745.83
Depreciation Added Back as non cash $         86,666.67 $       86,666.67 $         86,666.67
OCF EAT+Depreciation $     1,18,083.33 $    5,84,620.83 $    -2,85,079.17
Investment Given $ -5,20,000.00
CF $ -5,20,000.00   $ 1,18,083.33 $ 5,84,620.83 $ -2,85,079.17
  • Base cases figures are given
  • In best case, the Price and quantity will increase by 10% of base case scenario and VC and FC will go down by 10%
  • In worst case, the Price and quantity will fall by 10% of base case scenario and VC and FC will go up by 10%

Now we will calculate the NPV of each scenario:

Base case:

Year CF Discount Factor Discounted CF
0 $     -5,20,000.00 1/(1+0.1)^0= 1 1*-520000= $       -5,20,000.00
1 $       1,18,083.33 1/(1+0.1)^1= 0.909090909 0.909090909090909*118083.333333333= $        1,07,348.48
1 $       1,18,083.33 1/(1+0.1)^1= 0.909090909 0.909090909090909*118083.333333333= $        1,07,348.48
3 $       1,18,083.33 1/(1+0.1)^3= 0.751314801 0.751314800901578*118083.333333333= $            88,717.76
4 $       1,18,083.33 1/(1+0.1)^4= 0.683013455 0.683013455365071*118083.333333333= $            80,652.51
5 $       1,18,083.33 1/(1+0.1)^5= 0.620921323 0.620921323059155*118083.333333333= $            73,320.46
6 $       1,18,083.33 1/(1+0.1)^6= 0.56447393 0.564473930053777*118083.333333333= $            66,654.96
NPV = Sum of all Discounted CF $              4,042.65

Best case:

Year CF Discount Factor Discounted CF
0 $     -5,20,000.00 1/(1+0.1)^0= 1 1*-520000= $       -5,20,000.00
1 $       5,84,620.83 1/(1+0.1)^1= 0.909090909 0.909090909090909*584620.833333334= $        5,31,473.48
1 $       5,84,620.83 1/(1+0.1)^1= 0.909090909 0.909090909090909*584620.833333334= $        5,31,473.48
3 $       5,84,620.83 1/(1+0.1)^3= 0.751314801 0.751314800901578*584620.833333334= $        4,39,234.28
4 $       5,84,620.83 1/(1+0.1)^4= 0.683013455 0.683013455365071*584620.833333334= $        3,99,303.90
5 $       5,84,620.83 1/(1+0.1)^5= 0.620921323 0.620921323059155*584620.833333334= $        3,63,003.54
6 $       5,84,620.83 1/(1+0.1)^6= 0.56447393 0.564473930053777*584620.833333334= $        3,30,003.22
NPV = Sum of all Discounted CF $      20,74,491.91

Worst Case:

Year CF Discount Factor Discounted CF
0 $     -5,20,000.00 1/(1+0.1)^0= 1 1*-520000= $       -5,20,000.00
1 $     -2,85,079.17 1/(1+0.1)^1= 0.909090909 0.909090909090909*-285079.166666667= $       -2,59,162.88
1 $     -2,85,079.17 1/(1+0.1)^1= 0.909090909 0.909090909090909*-285079.166666667= $       -2,59,162.88
3 $     -2,85,079.17 1/(1+0.1)^3= 0.751314801 0.751314800901578*-285079.166666667= $       -2,14,184.20
4 $     -2,85,079.17 1/(1+0.1)^4= 0.683013455 0.683013455365071*-285079.166666667= $       -1,94,712.91
5 $     -2,85,079.17 1/(1+0.1)^5= 0.620921323 0.620921323059155*-285079.166666667= $       -1,77,011.73
6 $     -2,85,079.17 1/(1+0.1)^6= 0.56447393 0.564473930053777*-285079.166666667= $       -1,60,919.76
NPV = Sum of all Discounted CF $    -17,85,154.35

So the correct answer is none of the above